OpenStudy (anonymous):

Which expression is equivalent to (cos x)(tan(-x))? a) -sin x b) sin x c) -csc x d) csc x

OpenStudy (anonymous):

Firstly : \[\tan(-x) = -\tan(x)\]

OpenStudy (anonymous):

Secondly: \[\tan(x) = \frac{\sin(x)}{\cos(x)}\] Replace tan(x) with the above Identity, cancel out whatever is being cancelled and you have your answer with you then. :)